# Equations involving the modular $j$-function and its derivatives

### v1

@@ 1.1. Structure of the paper
We go over some basic preliminaries about the $j$ function and its derivatives. We also give the definition of Zariski density used in Theorem 1.2.
@@
We produce zero estimates for polynomials in $z$, $j$, $j^{\prime}$, $j^{\prime\prime}$ in order to determine when the quotients mentioned above have poles.
@@
First, in §7.1 we prove Zariski density for $\mathbf{j}$-homogeneous equations (Definition 7.1), which only involve $j(z)$, $j^{\prime}(z)$ and $j^{\prime\prime}(z)$, including the equation $j^{\prime\prime}(z)=0$. Finally, in §7.2 we prove Theorem 1.2 in full generality.
@@ 2. Preliminaries
Let $\operatorname{\mathbb{H}}$ denote the complex upper-half plane $\{z\in\mathbb{C}:\mathrm{Im}(z)>0\}$. The group $\mathrm{GL}_{2}^{+}(\operatorname{\mathbb{R}})$ of $2\times 2$ real invertible matrices with positive determinant acts on $\operatorname{\mathbb{H}}$ via linear fractional transformations:
@@
This action can be seen as a restriction of the action of $\mathrm{GL}_{2}(\operatorname{\mathbb{C}})$ on the Riemann sphere ${}\operatorname{\mathbb{C}}\cup\{\infty\}$. The modular group is defined as
@@
The modular $j$-function is defined as the unique holomorphic function $j:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ that satisfies
$j(\gamma z)=j(z)\text{ for every }\gamma\text{ in }\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})\text{ and every }z\text{ in }\operatorname{\mathbb{H}},$
and has a Fourier expansion (also known as a $q$-expansion) of the form
@@
In fact, $a_{k}\in\operatorname{\mathbb{Z}}$ for every $k\in\mathbb{N}$. The $j$-function induces an analytic isomorphism of Riemann surfaces $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})\backslash\operatorname{\mathbb{H}}\simeq\operatorname{\mathbb{C}}$.
Since $j$ is invariant under the action of $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$, we can study the behaviour of $j$ by looking at fundamental domains of the action of $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$ on $\operatorname{\mathbb{H}}$. The standard fundamental domain is the set
@@
We let $\overline{\mathbb{F}}$ denote the Euclidean closure of $\mathbb{F}$ (within the Riemann sphere). A diagram of the standard fundamental domain along with some of its $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$-translates is given in Figure 1, where $\rho:=\exp\left(\frac{2\pi i}{3}\right)$. When we refer to a fundamental domain, we will always mean a set of the form $\gamma\mathbb{F}$ for some $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$.
@@
It is well known ([Mah69]) that $j$ satisfies the following third-order differential equation (and none of lower order): [^2] Observe that the denominators in the equation correspond to the polynomials that must be omitted in Theorem 1.2.
This shows that the derivatives of $j$ of order at least 3 are rational over $j$, $j^{\prime}$ and $j^{\prime\prime}$.
A meromorphic function $f:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ is said to be $1$-periodic if $f(z+1)=f(z)$ for all $z\in\operatorname{\mathbb{H}}$. When $f$ is holomorphic and 1-periodic it has a $q$-expansion, that is to say, $f$ can be written as $f(z)=\sum_{k=-n}^{\infty}a_{k}q^{k}$, where $q=\exp(2\pi iz)$ and $a_{k}\in\operatorname{\mathbb{C}}$. In particular $j$, $j^{\prime}$ and $j^{\prime\prime}$ are all 1-periodic, and by differentiating (2.1) we can obtain the $q$-expansions of $j^{\prime}$ and $j^{\prime\prime}$:
@@
Observe that ${}\operatorname{\mathbb{Q}}\cup\{\infty\}$ forms a single orbit under the action of $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$. We call these elements the cusps of $j$. Given a fundamental domain $\gamma\mathbb{F}$, the cusp of $\gamma\mathbb{F}$ is the unique element of ${}\operatorname{\mathbb{Q}}\cup\{\infty\}$ contained in the Euclidean closure of $\gamma\mathbb{F}$ (where the closure is taken in the Riemann sphere).
Using (2.1) and the $q$-expansions of $j^{\prime}$ and $j^{\prime\prime}$, it is easy to see that for every $z\in\operatorname{\mathbb{H}}$ we have that each of the expressions $j(z+iy)$, $j^{\prime}(z+iy)$ and $j^{\prime\prime}(z+iy)$ grows exponentially to $\infty$ as $y\to+\infty$. In this case we sometimes write $z\to i\infty$ to emphasise that $z$ approaches $\infty$ by increasing its imaginary part, while the real part remains bounded. A similar behaviour takes place when $z$ approaches a rational number from within a fixed fundamental domain containing that rational number in its euclidean closure. We will give a precise description of this behaviour in §4.
@@
Let $F(X,Y_{0},Y_{1},Y_{2})$ be a polynomial over $\operatorname{\mathbb{C}}$. We say the equation $F(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0$ has a Zariski dense set of solutions if for any polynomial $G(X,Y_{0},Y_{1},Y_{2})$ which is not divisible by some irreducible factor of $F$, there is $z_{0}\in\operatorname{\mathbb{H}}$ such that $F(z_{0},j(z_{0}),j^{\prime}(z_{0}),j^{\prime\prime}(z_{0}))=0$ and $G(z_{0},j(z_{0}),j^{\prime}(z_{0}),j^{\prime\prime}(z_{0}))\neq 0$.
@@ 3. Existence of Solutions
We start with the Rouché method for proving the existence of solutions, but not yet their Zariski density.
@@
    If $u\in\operatorname{\mathbb{R}}\setminus\operatorname{\mathbb{Q}}$ and the sequence $\gamma_{k}=\begin{pmatrix}a_{k}&b_{k}\\ c_{k}&d_{k}\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$ is such that $\gamma_{k}z\to u$ as $k\to+\infty$, then $|c_{k}|\to+\infty$ as $k\to+\infty$.
@@
    Since every subsequence of $\gamma_{k}z$ tends to $u$, it suffices to show that $c_{k}$ is unbounded. Assume it is bounded, then we can choose a subsequence where the value of $c_{k}$ is constant. So assume $c_{k}=c$ is a constant sequence.
    Let $z=x+iy$. If $d_{k}$ is also bounded then we may assume it is constant. Then $a_{k}z+b_{k}=(a_{k}x+b_{k})+a_{k}yi$ must be convergent, hence $a_{k}$ must be convergent, and so constant. Then $b_{k}$ is also constant, for $a_{k}d_{k}-b_{k}c_{k}=1$, a contradiction.
@@
    Since $u=\frac{a}{c}$ with $\gcd(a,c)=1$, there are integers $m,l$ such that $am-cl=1$. Choose $b_{k}=l+ka,~{}d_{k}=m+kc$. Then
@@
Pick a point $u=\frac{a}{c}\in\operatorname{\mathbb{Q}}$, and choose a sequence $\gamma_{k}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ as in Lemma 3.1 such that $\gamma_{k}z\to u$ as $k\to+\infty$ for any $z\in\operatorname{\mathbb{H}}$. Let $B_{k}\coloneqq\gamma_{k}B$. By compactness of $B$, the function $r(\gamma_{k}z)$ tends to $r(u)$ uniformly for $z\in B$.
@@
We can now shrink $B$ to make sure that $H(u,\mathbf{j}(z))\neq 0$ on $\partial B$, so it is uniformly bounded away from 00 for $z\in\partial B$. In particular, $f(\gamma_{k}z)$ approaches infinity as $k\to+\infty$ uniformly for $z\in\partial B$. So for sufficiently large $k$ the inequality $|f(z)|>|r(z)|$ holds for all $z\in\partial B_{k}=\partial(\gamma_{k}B)=\gamma_{k}\partial B$, and we can apply Rouché’s theorem to these functions. Since $f$ has a zero in $B_{k}$, namely $\gamma_{k}\rho$, so does $f-r$. ∎
@@
Let $F(X,\mathbf{Y})\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]\setminus\operatorname{\mathbb{C}}[X]$. Let $U\subseteq\operatorname{\mathbb{C}}$ be an open set such that $U\cap\operatorname{\mathbb{R}}\neq\emptyset$ and $f:U\rightarrow\operatorname{\mathbb{C}}$ be a holomorphic function. Then the equation $F(z,\mathbf{j}(z))=f(z)$ has infinitely many solutions.
@@ 4. Solvability of Certain Equations Involving Periodic Functions
We recall that given a meromorphic function $f$ and a point $z_{0}$, the order of $f$ at $z_{0}$ is the unique integer $n$ such that $(z-z_{0})^{-n}f(z)$ is holomorphic and non-zero at $z_{0}$.
@@
Let $f_{0},\ldots,f_{n}:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ be $1$-periodic meromorphic functions and let $-\ell$ be the minimum order of $\frac{f_{k}}{f_{n}}$ at a fixed $z_{0}\in\operatorname{\mathbb{H}}$ for $k=0,\dots,n$.
If $\ell>0$, then for any sufficiently small disc $D$ centred at $z_{0}$ and for every sufficiently large $m\in\operatorname{\mathbb{Z}}$ the following equation has $\ell$ solutions, counted with multiplicity, in $m+(D\setminus\{z_{0}\})$
$f_{n}(z)z^{n}+f_{n-1}(z)z^{n-1}+\dots+f_{0}(z)=0.$
@@
By Rouché’s theorem, the number of zeroes of the functions $G(z+m)$ and $G(z+m)+H(z+m)={f_{0}(z)}^{-1}F(z+m)$ inside $D$ is the same. Since the former has a zero at $z_{0}$ of order $\ell$ and no other zero, the latter must also have $\ell$ zeroes in $D$, counted with multiplicity. Thus, ${f_{0}(z)}^{-1}F(z+m)$ has $\ell$ zeroes in $D$, and so $f_{0}^{-1}F$ has $\ell$ zeroes in $m+D$, counted with multiplicity.
Finally, note that $G(z_{0}+m)+H(z_{0}+m)=0$ holds for at most $n-1$ values of $m$, thus for $m$ sufficiently large, the above $\ell$ solutions in $m+D$ are actually in $m+(D\setminus\{z_{0}\})$.
@@
Following the notation of the proposition, when the $f_{k}$’s are polynomials in $j$, $j^{\prime}$, $j^{\prime\prime}$ and some $\frac{f_{k}}{f_{n}}$ has a pole in $\operatorname{\mathbb{H}}$, the above proposition applies. Instead, when the functions $\frac{f_{k}}{f_{n}}$ have no poles in $\operatorname{\mathbb{H}}$, we will rely on the asymptotic behaviour of $\frac{f_{k}}{f_{n}}$ towards the boundary of $\operatorname{\mathbb{H}}$. Specifically, we will prove an analogue of Proposition 4.1 in the case where $\frac{f_{k}(z)}{f_{n}(z)}\to\infty$ as $z\to i\infty$.
As the following example shows, we may also need to consider non-periodic functions which are asymptotically periodic. Dealing with those functions requires a considerably more sophisticated setup than the one in Proposition 1.5, so we first discuss the example in detail to clarify the choices made in the rest of this section.
@@
where $\alpha$ is to be determined later. In order to write this equation as a polynomial in $z$ with periodic coefficients, we apply the $z\mapsto-\frac{1}{z}$ transformation and get
@@
In this example the ratio of the coefficients of $z^{8}$ and $z^{10}$ actually has a pole at $i$. However, checking this in general requires sufficiently precise zero estimates at the conjugates of $\rho$ and $i$, which are hard to produce for polynomials involving $j^{\prime\prime}$ (see Example 6.7); on the other hand, we can provide sharp zero estimates for polynomials in $j$, $j^{\prime}$ only (see §6). The latter estimates turn out to be enough: for instance, when the original equation does not contain $z$, after the $z\mapsto-\frac{1}{z}$ transformation the coefficient of the lowest power of $z$ does not depend on $j^{\prime\prime}$. This is exemplified here by the coefficient of $z^{6}$. Hence our strategy hinges on the fact that the ratio between a particular coefficient not involving $j^{\prime\prime}$, which we can procure in all cases, and the leading coefficient has a pole or exponential growth on some fundamental domain.
In this particular example, the ratio in question is between the coefficients of $z^{6}$ and of $z^{10}$, thus the function
@@
We claim that $f(z)$ has no pole in $\operatorname{\mathbb{H}}$. Indeed, easy calculations show that the orders of the numerator at $\rho$ and $i$ (and their $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits) are equal to 6 and 3 respectively. The denominator has the same orders at these points, so $f$ has no poles. Moreover, choosing $\alpha=\frac{2}{\pi i}$ ensures the leading terms in the $q$-expansions of the two terms in the numerator cancel out. This then means that $f(z)$ tends to a constant as $z\to i\infty$. Therefore, Proposition 1.5 cannot be applied in this situation. However, $f(z)$ has exponential growth as we approach 00 from within a fundamental domain with a cusp at 00 (in fact, we shall prove that $f(z)$ must have exponential growth for most fundamental domains). Indeed, after applying the $z\mapsto-\frac{1}{z}$ transformation we get
@@
It is worth mentioning that for equations of the form $F(j,j^{\prime})=0$, where $F$ is a polynomial, the transformation $z\mapsto-\frac{1}{z}$ always guarantees that the ratio of the lowest and highest powers of $z$ has a pole in $\operatorname{\mathbb{H}}$ or exponential growth at $i\infty$. In general, Proposition 1.5 is sufficient when we deal with equations of the form $F(z,j(z),j^{\prime}(z))=0$, although the argument is somewhat more complicated.
@@
Let $\mathcal{P}$ denote the field of $1$-periodic meromorphic functions on $\operatorname{\mathbb{H}}$; recall that all such functions have a $q$-expansion, namely they can be written as meromorphic functions on the unit disc evaluated at $q=\exp(2\pi iw)$ for $w$ in $\operatorname{\mathbb{H}}$. We write $\mathcal{P}[w]$ and $\mathcal{P}(w)$ for the ring and the field generated by the variable $w$ over $\mathcal{P}$.
@@
It suffices to prove the conclusion for $f\in\mathcal{P}[w]$. Write $f(w)=\sum_{k=0}^{n}g_{k}(w)w^{n}$, with $g_{k}\in\mathcal{P}$.
Each $g_{k}(w)$ has a meromorphic $q$-expansion $\tilde{g}_{k}(q)$ on some neighbourhood of $q=0$. Let $e$ be the minimum order of $\tilde{g}_{k}(q)$ at $q=0$ for $k=0,\dots,n$. Then we can write the Laurent series
$G(w,q)=\sum_{k=0}^{n}\tilde{g}_{k}(q)w^{k}=\sum_{m=e}^{\infty}q^{m}p_{m}(w)$
where the coefficients $p_{m}(w)$ are polynomials of degree at most $n$. Let $d$ be the degree of $p_{e}(w)$ and $\alpha$ be the coefficient of $w^{d}$ in $p_{e}(w)$. It follows at once that in the region $|q|\leq|w|^{-(n+1)}$ we have
$G(w,q)=\alpha w^{d}q^{e}\left(1+O\left(w^{-1}\right)+O(q)\right).$
It now suffices to specialise at $q=\exp(2\pi iw)$ and observe that $|q|=e^{2\pi\operatorname{Im}w}\leq|w|^{-(n+1)}$ if and only if $\operatorname{Im}(w)\geq\frac{n+1}{2\pi}\log|w|$. ∎
This lemma allows us to make the following definition.
@@
Call order at $i\infty$ of $f\in\mathcal{P}(w)$, written $\operatorname{ord}_{w=i\infty}(f)$, the pair $(e,d)\in\operatorname{\mathbb{Z}}^{2}$ of integers such that for some $C\in\operatorname{\mathbb{R}}$, we have $f(w)\sim\alpha w^{-d}\exp(2\pi iew)$ for $w\to\infty$ in the region $\operatorname{Im}(w)\geq C\log|w|$.
@@
Here we consider $\operatorname{\mathbb{Z}}^{2}$ as an ordered group, ordered lexicographically. This makes $(\mathcal{P}(w),\operatorname{ord}_{w=i\infty})$ into a valued field.

### v3

@@ 1.1. Structure of the paper
We go over some basic preliminaries about the $j$-function and its derivatives. We also give the definition of Zariski density used in Theorem 1.2.
@@
We produce zero estimates for polynomials in $z$, $j$, $^{\prime}, j′′j^{\prime\prime}$ in order to determine when the quotients mentioned above have poles.
@@
First, in §7.1 we prove Zariski density for $\mathbf{j}$-homogeneous equations (Definition˜7.1), which only involve $j(z)$, $j^{\prime}(z)$ and $j^{\prime\prime}(z)$, including the equation $j^{\prime\prime}(z)=0$. Finally, in §7.2 we prove Theorem 1.2 in full generality.
@@ 2. Preliminaries
[…]
@@
[…]
@@
The modular $j$-function is defined as the unique $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$-automorphic function $j:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ satisfying $j(\rho)=0$ (recall that $\rho:=\exp\left(\frac{2\pi i}{3}\right)$; this notation will be kept throughout the paper), $j(i)=1728$ and $j(\infty)=\infty$ (this last condition should be understood as $\lim_{\mathrm{Im}(z)\to+\infty}j(z)=\infty$). In particular, this means that $j$ satisfies
[…]
and it is in particular 1-periodic, and by assumption meromorphic at $i\infty$. Its Fourier expansion (also known as a $q$-expansion) is of the form
@@
In fact, $a_{k}\in\operatorname{\mathbb{Z}}$ for every $k\in\mathbb{N}$. The $j$-function induces an analytic isomorphism of Riemann surfaces $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})\backslash\operatorname{\mathbb{H}}\simeq\operatorname{\mathbb{C}}$ (see [Lan87, Chapter 3, §3]).
[…]
@@
We let $\overline{\mathbb{F}}$ denote the Euclidean closure of $\mathbb{F}$ (within the Riemann sphere). A diagram of the standard fundamental domain along with some of its $\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$-translates is given in Figure˜1. When we refer to a fundamental domain, we will always mean a set of the form $\gamma\mathbb{F}$ for some $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$.
It is well known that $j$ satisfies the following third-order differential equation (and none of lower order, [Mah69]): [^2] Observe that the denominators in the equation correspond to the polynomials that must be omitted in Theorem 1.2.
@@
The functions $j$, $j^{\prime}$ and $j^{\prime\prime}$ are all 1-periodic and meromorphic at $i\infty$, and by differentiating (2.1) we can obtain the $q$-expansions of $j^{\prime}$ and $j^{\prime\prime}$:
@@
[…]
Using (2.1) and the $q$-expansions of $j^{\prime}$ and $j^{\prime\prime}$, it is easy to see that for every $x\in\operatorname{\mathbb{R}}$ we have that each of the expressions $j(x+iy)$, $j^{\prime}(x+iy)$ and $j^{\prime\prime}(x+iy)$ grows exponentially to $\infty$ as $y\to+\infty$. In this case we sometimes write $z\to i\infty$ to emphasise that $z$ approaches $\infty$ by increasing its imaginary part, while the real part remains bounded. A similar behaviour takes place when $z$ approaches a rational number from within a fixed fundamental domain containing that rational number in its euclidean closure. We will give a precise description of this behaviour in §4.
@@
[…]
$F(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0$
has a Zariski dense set of solutions if for any polynomial $G(X,Y_{0},Y_{1},Y_{2})$ which is not divisible by some irreducible factor of $F$, there is $z_{0}\in\operatorname{\mathbb{H}}$ such that
$F(z_{0},j(z_{0}),j^{\prime}(z_{0}),j^{\prime\prime}(z_{0}))=0\text{ and }G(z_{0},j(z_{0}),j^{\prime}(z_{0}),j^{\prime\prime}(z_{0}))\neq 0.$
@@ 3. Existence of Solutions
We start with the Rouché method for proving the existence of solutions, but not yet their Zariski density. We first recall the crucial theorem.
Theorem 3.1 (Rouché, see e.g. [Lan99, Chapter VI, §1, Theorem 1.6]).
Let $f,g$ be meromorphic functions on a complex domain $\Omega$. Let $C$ denote a simple closed curve which is homologous to 0 in $\Omega$ and such that $f$ has no zeroes or poles on $C$. If the inequality
$|g(z)|<|f(z)|$
holds for all $z$ on $C$, then the difference between the numbers of zeroes and poles in the interior of $C$ for the functions $f+g$ and $f$ is the same.
[…]
@@
    If $u\in\operatorname{\mathbb{R}}\setminus\operatorname{\mathbb{Q}}$ and the sequence $\gamma_{k}=\begin{pmatrix}a_{k}&b_{k}\\ c_{k}&d_{k}\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$ is such that $\gamma_{k}z\to u$ as $k\to+\infty$, then $|c_{k}|\to+\infty$ and $\frac{a_{k}}{c_{k}}\to u$ as $k\to+\infty$.
@@
    We first show that $|c_{k}|\to+\infty$. Since every subsequence of $\gamma_{k}z$ tends to $u$, it suffices to show that $c_{k}$ is unbounded. Assume it is bounded, then we can choose a subsequence where the value of $c_{k}$ is constant. So assume $c_{k}=c$ is a constant sequence.
    Conversely, assume now that $\gamma_{k}z\to u$. Let $z=x+iy$. If $d_{k}$ is also bounded then we may assume it is constant. Then $a_{k}z+b_{k}=(a_{k}x+b_{k})+a_{k}yi$ must be convergent, hence $a_{k}$ must be convergent, and so constant. Then $b_{k}$ is also constant, for $a_{k}d_{k}-b_{k}c_{k}=1$, a contradiction.
@@
    Since $|c_{k}|\to+\infty$, then using that
    $\frac{a_{k}z+b_{k}}{c_{k}z+d_{k}}\cdot\frac{c_{k}}{a_{k}}=\frac{a_{k}z+b_{k}}{a_{k}z+b_{k}+\frac{1}{c_{k}}}\to 1$
    we see that $\gamma_{k}z$ and $\frac{a_{k}}{c_{k}}$ must have the same limit.
@@
    Since $u=\frac{a}{c}$ with $\gcd(a,c)=1$, there are integers $m,l$ such that amcl=1am-cl=1. Choose $b_{k}=l+ka,~d_{k}=m+kc$. Then
@@
Pick a point $u=\frac{a}{c}\in\operatorname{\mathbb{Q}}$, and choose a sequence $\gamma_{k}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ as in Lemma˜3.2((ii)) such that $\gamma_{k}z\to u$ as $k\to+\infty$ for any $z\in\operatorname{\mathbb{H}}$. Let $B_{k}\coloneqq\gamma_{k}B$. By compactness of $B$, the function $r(\gamma_{k}z)$ tends to $r(u)$ uniformly for $z\in B$.
@@
We can now shrink $B$ to make sure that $H(u,\mathbf{j}(z))\neq 0$ on $\partial B$, so it is uniformly bounded away from 0 for $z\in\partial B$. In particular, $f(\gamma_{k}z)$ approaches infinity as $k\to+\infty$ uniformly for $z\in\partial B$. So for sufficiently large $k$ the inequality $|f(z)|>|r(z)|$ holds for all $z\in\partial B_{k}=\partial(\gamma_{k}B)=\gamma_{k}\partial B$, and we can apply Rouché’s theorem to these functions. Since $f$ has a zero in $B_{k}$, namely $\gamma_{k}\rho$, so does frf-r. ∎
@@
Let $F(X,\mathbf{Y})\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]\setminus\operatorname{\mathbb{C}}[X]$. Let $U\subseteq\operatorname{\mathbb{C}}$ be an open set such that $U\cap\operatorname{\mathbb{R}}\neq\emptyset$ and let $f:U\rightarrow\operatorname{\mathbb{C}}$ be a holomorphic function. Then the equation $F(z,\mathbf{j}(z))=f(z)$ has infinitely many solutions.
@@ 4. Solvability of Certain Equations Involving Periodic Functions
We recall that given a meromorphic function $f$, not identically 0, and a point $z_{0}$, the order of $f$ at $z_{0}$ is the unique integer $n$ such that $(z-z_{0})^{-n}f(z)$ is holomorphic and non-zero at $z_{0}$.
@@
Let $f_{0},\ldots,f_{n}:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ be $1$-periodic meromorphic functions and let -$\ell$ be the minimum order of $\frac{f_{k}}{f_{n}}$ at a fixed $z_{0}\in\operatorname{\mathbb{H}}$ for $k=0,\dots,n-1$.
If $\ell>0$, then for any sufficiently small disc $D$ centred at $z_{0}$ and for every sufficiently large $m\in\operatorname{\mathbb{Z}}$, the equation
$f_{n}(z)z^{n}+f_{n-1}(z)z^{n-1}+\dots+f_{0}(z)=0$
has $\ell$ solutions, counted with multiplicity, in $m+(D\setminus\{z_{0}\})$.
@@
By Rouché’s Theorem˜3.1, the number of zeroes of the functions $G(z+m)$ and $G(z+m)+H(z+m)={f_{0}(z)}^{-1}F(z+m)$ inside $D$ is the same. Since the former has a zero at $z_{0}$ of order $\ell$ and no other zero, the latter must also have $\ell$ zeroes in $D$, counted with multiplicity. Thus, ${f_{0}(z)}^{-1}F(z+m)$ has $\ell$ zeroes in $D$, and so $f_{0}^{-1}F$ has $\ell$ zeroes in $m+D$, counted with multiplicity.
Finally, note that $G(z_{0}+m)+H(z_{0}+m)=0$ holds for at most n1n-1 values of $m$, thus for $m$ sufficiently large, the above $\ell$ solutions in $m+D$ are actually in $m+(D\setminus\{z_{0}\})$.
@@
[…]
As the following example shows, we may also need to consider non-periodic functions which are asymptotically periodic. Dealing with those functions requires a considerably more sophisticated setup than the one in Proposition˜1.6, so we first discuss the example in detail to clarify the choices made in the rest of this section.
@@
where $\alpha$ is to be determined later. In order to write this equation as a polynomial in $z$ with periodic coefficients, we apply the $z\mapsto-\frac{1}{z}$ transformation [^3] The action of the group $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ is generated by the transformations $z\mapsto z+1$ and $z\mapsto-\frac{1}{z}$. Since our functions are invariant under the former, it is natural to apply the latter; while $j$ is invariant under it, $j^{\prime}$ and $j^{\prime\prime}$ are not, and we take advantage of this fact. and get
@@
In this example the ratio of the coefficients of $z^{8}$ and $z^{10}$ actually has a pole at $i$. However, checking for poles among such ratios in a general equation requires sufficiently precise zero estimates at the conjugates of $\rho$ and $i$, which are hard to produce for polynomials involving $j^{\prime\prime}$ (see Example˜6.7); on the other hand, we can provide sharp zero estimates for polynomials in $j$, $j^{\prime}$ only (see §6). The latter estimates turn out to be enough: for instance, when the original equation does not contain $z$, after the $z\mapsto-\frac{1}{z}$ transformation the coefficient of the lowest power of $z$ does not depend on $j^{\prime\prime}$. This is exemplified here by the coefficient of $z^{6}$. Hence our strategy hinges on the fact that the ratio between a particular coefficient not involving $j^{\prime\prime}$, which we can procure in all cases, and the leading coefficient has a pole or exponential growth in some fundamental domain.
In this particular example, the ratio in question is between the coefficients of $z^{6}$ and $z^{10}$, thus the function
@@
[…]
@@
[…]
@@
[…]
Also, given an unbounded region $U\subseteq\operatorname{\mathbb{C}}$ and two meromorphic functions $f,g$ on $U$, we write $f\sim g$ for $w\to\infty$ in $U$ to mean that the limit of the ratio $\frac{f(w)}{g(w)}$ tends to 1 as $w$ approaches infinity from within $U$.
@@
It suffices to prove the conclusion for $f\in\mathcal{P}[w]$. Write $f(w)=\sum_{k=0}^{n}g_{k}(w)w^{k}$, with $g_{k}\in\mathcal{P}$.
Each $g_{k}(w)$ has a meromorphic $q$-expansion $\tilde{g}_{k}(q)$ which converges on some neighbourhood of $q=0$. Let $e$ be the minimum order of $\tilde{g}_{k}(q)$ at $q=0$ for $k=0,\dots,n$. Let $\alpha_{k}\in\operatorname{\mathbb{C}}$ be such that $\tilde{g}_{k}(q)=q^{e}(\alpha_{k}+O(q))$. Let $d$ be the maximum $k$ such that $\tilde{g}_{k}(q)$ has order $e$ at $q=0$, namely such that $\alpha_{k}\neq 0$. In the region $|q|\leq|w|^{-(n+1)}$ we have $O(q)=O(w^{-(n+1)})$, in which case
$G(w,q)=\sum_{k=0}^{n}\tilde{g}_{k}(q)w^{k}=q^{e}w^{d}\sum_{k=0}^{n}\left(\alpha_{k}+O\left(w^{-(n+1)}\right)\right)w^{k-d}=\alpha_{d}q^{e}w^{d}\left(1+O\left(w^{-1}\right)\right).$
It now suffices to specialise at $q=\exp(2\pi iw)$ and observe that $|q|=e^{-2\pi\operatorname{Im}w}\leq|w|^{-(n+1)}$ if and only if $\operatorname{Im}(w)\geq\frac{n+1}{2\pi}\log|w|$. ∎
It follows at once that all functions in $\mathcal{P}(w)$ are ‘asymptotically periodic’ in the sense that $f(w+1)\sim f(w)$ for $w\to\infty$ in the above region. Moreover, the lemma allows us to make the following definition.
@@
Call the order at $i\infty$ of $f\in\mathcal{P}(w)$, written $\operatorname{ord}_{w=i\infty}(f)$, the pair $(e,d)\in\operatorname{\mathbb{Z}}^{2}$ of integers such that for some $\alpha\in\operatorname{\mathbb{C}}$ and some $C\in\operatorname{\mathbb{R}}$, we have $f(w)\sim\alpha w^{-d}\exp(2\pi iew)$ for $w\to\infty$ in the region $\operatorname{Im}(w)\geq C\log|w|$.
@@
Here we consider $\operatorname{\mathbb{Z}}^{2}$ as an ordered group, ordered lexicographically. This makes $(\mathcal{P}(w),\operatorname{ord}_{w=i\infty})$ into a valued field, and we have for instance that $f(w)\to\infty$ in a suitable region $\operatorname{Im}(w)\geq C\log|w|$ if and only if $\operatorname{ord}_{w=i\infty}f(w)<(0,0)$.
We can now set up a generalisation of Proposition˜4.1 that will cover our application to $j$. Let us fix the following data:
  • a polynomial $F(z,w)=\sum_{k=0}^{n}z^{k}f_{k}(w)$ where each $f_{k}$ is in $\mathcal{P}(w)$ and $f_{n}\neq 0$;
  • a value $s$ which is either 0 or $1$.
We look for the zeroes of functions of the form $F_{r}(w)\coloneqq F(r+sw,w)$ for $r$ varying among the real numbers. For each $r$ sufficiently large, we pick a suitable rectangle $\Xi_{r}$, pictured in Figure˜2 and defined in Proposition˜4.8, and integrate the logarithmic derivative $\frac{F_{r}^{\prime}}{F_{r}}$ along the boundary of $\Xi_{r}$. Provided that $F_{r}$ does neither have zeroes nor poles on such a boundary, by the Argument Principle (see Theorem˜4.7), the value of the integral counts the difference between the number of zeroes and poles, with multiplicity, inside $\Xi_{r}$. We will choose $\Xi_{r}$ so that the integral has positive value.
We first parametrise the roots of $F(z,w)$ as a polynomial in $z$ in terms of $w$ varying in a suitable region. The resulting functions, which by construction are algebraic over $\mathcal{P}(w)$, admit an order at $i\infty$ which may be a pair of rational numbers, rather than only integers.